OpenStudy (anonymous):

A particle undergoes two displacements, measured from the positive x-axis, with counterclockwise positive. The first has a magnitude of 11 m and makes an angle of 69 ◦ with the positive x axis. The resultant displacement has a magnitude of 10 m directed at an angle of 120 ◦ from the positive x axis. Find the magnitude of the second displacement.

OpenStudy (anonymous):

I know I need to find B=\[\sqrt{bx^2+by^2}\] and Bx=Cx-Ax and By=Cy-Ay

OpenStudy (anonymous):

help plz??

OpenStudy (anonymous):

break both the resultant and the first displacement into components of x and y. Subtract them into the components of displacement B. Use the Pythagorean theorem and you should get: <b>=9.09m use tangent fot the angle to get: angle=maybe 7.4 degrees

OpenStudy (anonymous):

magnitude of the second displacement (b)= b^2=11*11+10*10-10*11*cos(51) b^2=221-81.6=139.36 b=11.8